OpenStudy (anonymous):

Simplify. Remember to use absolute-value notation when necessary. If a root cannot be simplified, state this. 12square root (-10)^12=

OpenStudy (anonymous):

\[\Large 12\sqrt{-10^{12}} \quad ???\]

OpenStudy (anonymous):

or ...\[\Large \sqrt[12]{-10^{12}}\]

OpenStudy (anonymous):

ahhh... I mean in bracket like thiss... \[\Large 12\sqrt{(-10)^{12}}\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yes. Now I don't understand how to simplify this

OpenStudy (anonymous):

Yes which one ?? .. (I'm kreshnik :D\[\Large 12\sqrt{(-10)^{12}} \quad ???\])

OpenStudy (anonymous):

The last one you just posted

OpenStudy (anonymous):

... Remember that ... you can rewrite it like..\[\Large 12\sqrt{(-10)^{12}}=12\sqrt{10^{12}}=12\cdot 10^{\frac{12}{2}}\] Now you have...\[\Large 12\cdot 10^{6}\] Use calculator if you want an exact number !!

OpenStudy (inkyvoyd):

12 million

OpenStudy (anonymous):

whoaa... hahaha LOL

OpenStudy (anonymous):

Now I'm kreshnik xD :P